# A review of Bayesian sensor-based estimation and uncertainty quantification of aerodynamic flows

### v2

@@ 3.4.1 One true vortex and one-vortex estimator
Figure 8 illustrates the estimation results for the trajectory and strength of the vortex passing over a cylinder in uniform flow. The spatiotemporal observation map in Fig. 8(a) depicts the surface pressure recorded by all of the sensors over time. The surface pressure sensors positioned on the left side of the cylinder initially respond to the vortex’s approach around $t=t_{s}$, when the vortex is closest to these sensors (see the location of the vortex at this moment in Fig. 8(b)). As the vortex moves along the upper surface, sensors in its immediate vicinity become increasingly responsive. An ensemble of vortices is initially placed in the covariance ellipse illustrated in Fig. 8(b), with the true initial vortex location shown for reference. The estimated trajectory of the vortex converges to the true trajectory after only three time steps and remains close to this trajectory with consistently low uncertainty set by the covariance inflation. The comparison of the ensemble mean’s history with the true trajectory is depicted in Fig. 9 for the vortex coordinates and strength. The plots show the immediate effect of using measurements to update the state from its initial prior ensemble.
@@
Figure 8: One true vortex and one-vortex estimator. (a) The spatiotemporal plot of sensor pressures, with sensor indices starting from the rightmost point on the cylinder’s surface and progressing counterclockwise (refer to Fig. 7). (b) True vortex trajectory (black solid line) alongside the estimated trajectory (yellow dotted line). The initial ensemble mean and covariance ellipse are shown. Gray squares on the cylinder represent sensors.
@@
Figure 10 presents the first three eigenvectors of the measurement-space Gramian at three different time steps as the vortex passes by the cylindrical body. These modes demonstrate the dominant ways in which measurement data is aggregated to update the state estimate. The figure shows that when the vortex is far from the body, sensor readings contribute in a relatively broadly over the whole cylinder, suggesting that distance localization would be a poor strategy for improving the performance of the sEnKF. As the vortex nears the cylinder, measurements from sensors in its vicinity become more responsive to flow changes, and these dominant modes only gather information from a few sensors. Higher modes generally have more complex features and less spatial compactness.
@@ 3.4.2 Five true vortices and five-vortex estimator
As in the static inference case, the complexity of the problem increases significantly when multiple vortices are present in the flow. To demonstrate the capability of the LREnKF in handling (and illuminating) such complex estimation scenarios, we set up a case in which the true flow consists of five point vortices initially placed upstream of the cylinder in an arbitrary configuration. We use the LREnKF to estimate the trajectories and strengths of all five vortices from the measurements obtained from the array of surface pressure sensors. The true initial positions and strengths of the vortices are given by $(x^{*}_{0}/R_{c},y^{*}_{0}/R_{c},\Gamma/UR_{c})=(-3,-0.8,-1)$, $(-2.8,-0.5,-1.1)$, $(-2.9,0,1.2)$, $(-2.7,0.4,1.3)$ and $(-2.8,0.8,1.4)$. An ensemble of $M=50$ members is used to estimate the $n=15$ components of the state of the five vortices. The initial ensemble is drawn from a Gaussian prior with a mean that is offset from the true value, similar to the single vortex case. To offset the relabeling symmetry discussed in Section 3, the Gaussians associated with each vortex do not overlap with one another. The offsets are randomly drawn from uniform distributions $(x_{\text{offset}}\sim\mathcal{U}(0,0.4)$, $y_{\text{offset}}=0$, and $\Gamma_{\text{offset}}\sim\mathcal{U}(-0.4,0.4))$, where $\mathcal{U}$ denotes a uniform distribution, and the initial prior’s standard deviations for each vortex are $\sigma_{x}=0.2$, $\sigma_{y}=0.1$ and $\sigma_{\Gamma}=0.1$.
@@
Figure 11: Five true vortices and five vortex estimation. (a) The spatiotemporal plot of surface sensor pressures, with sensor indices starting from the rightmost point on the cylinder’s surface and progressing counterclockwise (refer to Fig. 7). (b) Evolution of the ranks of the state and measurement Gramians. (c) True vortex trajectories (solid lines) and the predicted trajectories (dashed lines), each accompanied by $95\%$ confidence intervals. The initial ensemble means, along with initial covariance ellipses, are shown. Gray squares on the cylinder represent sensors. (d) The true (solid lines) and predicted (dashed lines) circulations for vortices 1 and 5 (with colors consistent with panel (c)), with $95\%$ confidence intervals.
Figure 11(a) presents the transient response of the pressure sensors during the vortex encounter. The observation map clearly indicates that sensors at the front of the cylinder, progressed toward the upper and lower sections, exhibit heightened responses as the vortices traverse the cylinder during the time interval $t\in[t_{s1},t_{s2}]$. For a detailed view of the true vortex positions at these specific instants, refer to Figure 11(c). This panel also shows the predicted trajectories of the five vortices, and Figure 11(d) presents the predicted strengths of two of the vortices, all with associated $95\%$ confidence intervals. The initial ensembles are shown in panel (c) with their mean and covariance ellipses shown in red. The trajectories demonstrate the complex interactions among the vortices and with the body. Initially, with the vortices far from the body, the estimator predicts a high-uncertainty region around each trajectory, encompassing the true state. As the vortices approach the body, the estimator’s predictions generally become more accurate and uncertainty drops significantly. Uncertainty also tends to grow during intervals in which vortices come close to one another, and consequently, their forecast motions are most sensitive to small perturbations.
@@
Figure 12: Five true vortices and five vortex estimation: (a) 1, 2, 3, and 10th eigenvectors of observation Gramian $C_{\mathsf{y}}$, $(u_{1},u_{2},u_{3},u_{10})$ are depicted at four distinct time instants as the vortices (indicated by black circles) advect with the freestream flow past a cylinder, (b) the evolution of 1, 2, 3, and 10th eigenvalues of the observation Gramian.
@@ 3.4.3 Five true vortices and three-vortex estimator
In all other aspects, the problem remains unchanged from the previous example, except the initial true states are slightly different, given by $(x^{*}_{0}/R_{c},y^{*}_{0}/R_{c},\Gamma/UR_{c})=(-3,-0.8,-1)$, $(-2.8,-0.5,-1.1)$, $(-2.9,0,1.2)$, $(-2.7,0.5,1.3)$ and $(-2.8,0.8,1.4)$. Additionally, the dimension of the estimated state is now reduced to $n=9$, compared to 15 in the previous case. Figure 13(a) shows the true and estimated trajectories. Obviously the three estimated vortices cannot precisely follow the trajectories of the five true ones. However, they do appear to estimate the cluster trajectories well. Indeed, Fig. 13(b) confirms this clustering effect in the estimated vortex strengths: the estimated vortex nearest the two negatively signed true vortices below the cylinder is estimated to have a strength nearly equal to the sum of those two true vortices, denoted by $\Gamma_{1}$ and $\Gamma_{2}$. Similarly, the green and blue estimated vortices, which are positioned between the uppermost true vortices, $\Gamma_{3}$ through $\Gamma_{5}$, approximate the combined circulation of these three vortices.
@@
Figure 13: Five true vortices and three vortex estimation. (a) True vortex trajectories alongside the predicted trajectories, each accompanied by $95\%$ confidence intervals. The initial ensemble means and covariance ellipses are shown. Gray squares on the cylinder represent sensors. (b) True and estimated and vortex circulations (with colors consistent with panel (a)), with $95\%$ confidence intervals.
@@ 3.5 Comments on sequential estimation
In Section 2 we stressed the importance of a good prior for successful flow estimation. In sequential estimation we achieve this by leveraging the knowledge obtained in previous steps. To underscore this, we return to the example of da Silva and Colonius [23], in which they sought to estimate the grid-based velocity field of flows past various bodies with a modest number of sensors. Without a good prior, it is hopeless for the measurements from a few velocity sensors to correct the high-dimensional grid velocity. However, the forecast model (the Navier–Stokes equations and boundary conditions) is deterministic, so even if the initial ensemble is drawn from a poor distribution—e.g., a spatially-uncorrelated Gaussian distribution, or white noise—the state’s ensemble is shaped over several forecast steps into a narrower and physically-plausible part of the state space. Indeed, in some of the examples in [23], the Navier-Stokes updates move the ensemble toward a trajectory on which the only uncertainty is in the phase of vortex shedding. Provided that there is sufficient initial variance of this phase among the ensemble members, the measurements are readily able to synchronize the evolving state onto the true phase. Even so, the benefits of this sequential improvement are still assisted by a good initial prior. Da Silva and Colonius [23] found the best performance when the initial states in their ensemble of $M$ members were drawn from the space spanned by the leading $M$ proper orthogonal decomposition (POD) modes of the true flow. In the vortex estimation examples in this section, our initial prior consisted of non-overlapping Gaussian distributions for each estimator vortex, each of which was reasonably close to a true vortex.
@@
The last example in Section 3.4, in which we interpreted measurements from a higher-dimensional system (five vortices) with a lower-dimensional representation (three vortices), provides an opportunity to make an important and more general point. The estimated trajectories in Fig. 13 show that the state covariance is quite small during most of the vortex–body interaction. However, it is crucial that we do not interpret this small covariance as precise knowledge of the true state. Rather, it represents a precise interpretation of the given measurements in the space in which we have confined our search. But that three-vortex space is a lower-dimensional manifold within the true space, and is thus inherently accompanied by a deterministic model error that we have no hope of estimating with our simplified state representation and observation model [28]. Fluid dynamics is replete with approximate representations of the flow field, each of which has a companion model error due to the restriction onto the simpler manifold, e.g., truncation error in grid-based CFD representations, subgrid-scale error in turbulence models, residual errors in reduced-order modal decompositions. For example, in the aforementioned study of da Silva and Colonius [23], in which they assembled their initial flow states from random combinations of POD modes, they might have instead chosen to work exclusively in the state space of POD coefficients. Some researchers have shown that it is possible to augment the state vector with the error itself and, by virtue of a model that relates it to the measurements, estimate this error along with the simplified flow state [29].
@@
Finally, we note that the LREnKF has provided a principled way to reduce the rank of the filtering so that the effect of spurious errors is mitigated, but there are a variety of other methods that improve the ensemble Kalman filter and we have only briefly listed a few. We also note that one could generalize the linear Kalman update (32) to a nonlinear form, thereby relaxing the assumption of Gaussian distributions in the analysis step. Such is the goal of the nonlinear ensemble filter of Spantini et al. [30]. Recall that the joint distribution $\pi(\mathsf{x},\mathsf{y})$ can be partitioned in two ways in equation (5): into the prior and likelihood, or into the marginal $\pi(\mathsf{y})$ and the conditional $\pi(\mathsf{x}|\mathsf{y})$ (i.e., the posterior, which we ultimately seek). Figure 1 illustrated these partitions for a linear observation operator, and we made use of the nice properties of the Gaussian to extract the posterior. However, another way to extract this posterior is to map the joint distribution (the ellipse in Figure 1) into a standard normal distribution $\mathcal{N}(0,I_{n+d})$ (a sphere of radius 1). The latter is trivial to partition into $\mathcal{N}(0,I_{d})\mathcal{N}(0,I_{n})$, and we can map the latter $n$-dimensional part back to the state space for the desired posterior. In fact, this invertible mapping is always possible for any distribution, and this is the basic idea used in [30]. For this to be useful, we first have to have some knowledge of the joint distribution $\pi(\mathsf{x},\mathsf{y})$, and we can obtain such knowledge by assembling an ensemble of prior states with their associated predicted noisy measurements (8). The method has been preliminarily evaluated in an aerodynamics problem by Le Provost et al. [31], but further investigation is needed to explore its merits in this context.
@@ 4 Neural networks and other learned mappings
Deep neural networks (NN) and other tools from machine learning have provided a fertile ground in recent years for learning operators that estimate flow quantities from aerodynamic measurements. Zhong et al. [32] have provided a pertinent example of this in the unsteady aerodynamics context. In that work, a small number of surface pressure measurements obtained from a CFD simulation of an airfoil in a disturbed flow were provided to a multi-layer perceptron (MLP) network. The network’s resulting vector was a high-dimensional representation of the flow field (though still lower dimensional than the flow itself). This vector was then used by a convolutional neural network (CNN) to reconstruct the vorticity field or by another MLP to estimate the instantaneous lift. Morra et al. [33] have recently shown that a trained deep operator network can be used in place of the full compressible Navier–Stokes equations to map upstream perturbations to downstream surface pressure measurements in a high-speed transitional boundary layer on a cone. They successfully used this approximate operator to estimate the amplitude of an upstream disturbance from downstream measurements taken from an experiment of the same geometry. Our objective in this paper is not to review the tools from machine learning and their application to fluid dynamics. Rather, our goal is to examine how to use these tools to support the stochastic flow estimation tasks that form the theme of this paper.
Among current architectures of deep neural networks, the autoencoder provides the most powerful for distilling data into low-dimensional representations. During unsupervised training on a set of high-dimensional vectors, an autoencoder simultaneously learns to compress this data into a low-dimensional latent space and to expand latent vectors back to the high-dimensional space with minimal error. We specifically call attention to the lift-augmented autoencoder developed recently by Fukami and Taira [34] (and illustrated in Figure 14), in which the addition of a NN to predict lift helps to regularize an autoencoder that learns to compress vorticity field snapshots from a wide variety of undisturbed and disturbed two-dimensional flows past an airfoil at several angles of attack. The resulting latent space is three-dimensional, i.e., only three components are needed to uniquely specify the instantaneous flow field!
Figure 14: Lift-augmented autoencoder for non-linear reduced order of flow states. Adapted from [34].
@@
As with physics-based operators, our overarching objective with learned operators is to understand how uncertainty in the measurements $\mathsf{y}$ translates into uncertainty in our state estimate $\mathsf{x}$—in this case, via uncertainty in the latent state $\boldsymbol{\xi}$. If we combine the NN operators with models for the measurement and process noise, e.g., zero-mean Gaussians, then we have the ingredients of the sequential filtering equations—the transition probability (27) and the likelihood (29)—and these enable us to track uncertainty through the flow prediction. But there are important questions to explore before pursuing this: How does uncertainty propagate through the latent space? And how does uncertainty in the model itself enter the problem? In this section we will explore these questions by focusing on the approximate inverse observation operator, $\tilde{g}(\mathsf{y};w)$, which maps measurements $\mathsf{y}$ to the latent state $\boldsymbol{\xi}$. We will use this opportunity to quantify the various forms of uncertainty in this mapping.
Figure 15: Neural network architecture for flow state prediction from sparse noisy pressure measurements $\boldsymbol{p}$, adapted from [35]. Measurement noise is translated to the predicted latent vector by learning the parameters of a multivariate Gaussian distribution in the latent space. The right panel depicts the low-dimensional latent space representation of disturbed flow fields. Colored trajectories correspond to baseline (undisturbed) cases at five different angles of attack, while gray trajectories represent disturbed cases, with the highlighted black trajectory emphasizing a specific disturbed scenario.
In our aerodynamics context, a representative example of this inverse mapping was investigated in [35], in which a lift-augmented autoencoder was trained on CFD simulation data from the same undisturbed and disturbed airfoil flows as in [34]. After training the autoencoder, the three-dimensional latent vector $\boldsymbol{\xi}_{j}$ for each training snapshot $n$ was paired with a measurement vector, $\mathsf{y}_{j}$, of the corresponding CFD-computed pressures at eleven evenly-spaced sensors on the surface of the airfoil. This latter vector also included the sensors’ position coordinates to distinguish angles of attack and to respect the sensors’ relative positioning. We will denote the training set of all such data pairs by $D=\{\mathsf{y}_{j},\boldsymbol{\xi}_{j}\}_{j=1}^{M}$.
@@
However, because the neural network provides us with access to the Jacobian $\nabla_{\mathsf{y}}\tilde{g}$ (through automatic differentiation), we can use this to characterize the uncertainty propagation more accurately. One such analytical method, stable distribution propagation, was introduced by Petersen et al. [36] and is applicable to Gaussian and Cauchy-distributed uncertainties. The core idea is to approximate the network’s behavior via local linearization. In the context of $\tilde{g}$, the expected value and variance of the predicted output are estimated as

### v3

@@ 3.4.1 One true vortex and one-vortex estimator
Figure 8 illustrates the estimation results for the trajectory and strength of the vortex passing over a cylinder in uniform flow. The spatiotemporal plot in Fig. 8(a) depicts the surface pressure recorded by all of the sensors over time. The surface pressure sensors positioned on the left side of the cylinder initially respond to the vortex’s approach around $t=t_{s}$, when the vortex is closest to these sensors (see the location of the vortex at this moment in Fig. 8(b)). As the vortex moves along the upper surface, sensors in its immediate vicinity become increasingly responsive. An ensemble of vortices is initially placed in the covariance ellipse illustrated in Fig. 8(b), with the true initial vortex location shown for reference. The estimated trajectory of the vortex converges to the true trajectory after only three time steps and remains close to this trajectory with consistently low uncertainty set by the covariance inflation. The comparison of the ensemble mean’s history with the true trajectory is depicted in Fig. 9 for the vortex coordinates and strength. The plots show the immediate effect of using measurements to update the state from its initial prior ensemble.
@@
Figure 8: One true vortex and one-vortex estimator. (a) Surface pressure sensor histories (indexed as in Fig. 7). (b) Vortex trajectories: true (black solid line) and estimated (yellow dotted line). Initial ensemble mean and covariance ellipse are shown. Surface sensors shown as gray squares.
@@
Figure 10 presents the first three eigenvectors of the measurement-space Gramian at three different time steps as the vortex passes by the cylindrical body. These modes demonstrate the dominant ways in which measurement data is aggregated to update the state estimate. The figure shows that when the vortex is far from the body, sensor readings contribute relatively broadly over the whole cylinder, suggesting that distance localization would be a poor strategy for improving the performance of the sEnKF. As the vortex nears the cylinder, measurements from sensors in its vicinity become more responsive to flow changes, and these dominant modes only gather information from a few sensors. Higher modes generally have more complex features and less spatial compactness.
@@ 3.4.2 Five true vortices and five-vortex estimator
As in the static inference case, the complexity of the problem increases significantly when multiple vortices are present in the flow. To demonstrate the capability of the LREnKF in handling (and illuminating) such complex estimation scenarios, we set up a case in which the true flow consists of five point vortices initially placed upstream of the cylinder in an arbitrary configuration. We use the LREnKF to estimate the trajectories and strengths of all five vortices from the measurements obtained from the array of surface pressure sensors. The true initial positions and strengths of the vortices are given by $(x^{*}_{0}/R_{c},y^{*}_{0}/R_{c},\Gamma/UR_{c})=(-3,-0.8,-1)$, (2.8,0.5,1.1)(-2.8,-0.5,-1.1), (2.9,0,1.2)(-2.9,0,1.2), (2.7,0.4,1.3)(-2.7,0.4,1.3) and (2.8,0.8,1.4)(-2.8,0.8,1.4). An ensemble of $M=50$ members is used to estimate the $n=15$ components of the state of the five vortices. The initial ensemble is drawn from a Gaussian prior with a mean that is offset from the true value, similar to the single vortex case. To offset the relabeling symmetry discussed in Section 3, the Gaussians associated with each vortex do not overlap with one another. The offsets are randomly drawn from uniform distributions $(x_{\text{offset}}\sim\mathcal{U}(0,0.4)$, $y_{\text{offset}}=0$, and $\Gamma_{\text{offset}}\sim\mathcal{U}(-0.4,0.4))$, where $\mathcal{U}$ denotes a uniform distribution, and the initial prior’s standard deviations for each vortex are $\sigma_{x}=0.2$, $\sigma_{y}=0.1$ and $\sigma_{\Gamma}=0.1$.
@@
Figure 11: Five true vortices and five-vortex estimator. (a) Surface pressure sensor histories. (b) Ranks of state and measurement Gramians. (c) Vortex trajectories, as in Fig. 8. (d) True (solid lines) and predicted (dashed lines) circulations for vortices 1 and 5 (colors consistent with panel (c)), with $95\%$ confidence intervals.
Figure 11(a) presents a spatiotemporal plot of the transient response of the surface pressure sensors during the vortex encounter. This plot clearly indicates that sensors at the front of the cylinder, and then those along the upper and lower sections, exhibit heightened responses as the vortices traverse the cylinder during the time interval $t\in[t_{s1},t_{s2}]$. For a detailed view of the true vortex positions at these specific instants, refer to Figure 11(c). This panel also shows the predicted trajectories of the five vortices, and Figure 11(d) presents the predicted strengths of two of the vortices, all with associated $95\%$ confidence intervals. The initial ensembles are shown in panel (c) with their mean and covariance ellipses shown in red. The trajectories demonstrate the complex interactions among the vortices and with the body. Initially, with the vortices far from the body, the estimator predicts a high-uncertainty region around each trajectory, encompassing the true state. As the vortices approach the body, the estimator’s predictions generally become more accurate and uncertainty drops significantly. Uncertainty also tends to grow during intervals in which vortices come close to one another, and consequently, their forecast motions are most sensitive to small perturbations.
@@
Figure 12: Five true vortices and five-vortex estimator: (a) 1, 2, 3, and 10th eigenvectors of observation Gramian $C_{\mathsf{y}}$, $(u_{1},u_{2},u_{3},u_{10})$ are depicted at four distinct time instants as the vortices (indicated by black circles), (b) the evolution of 1, 2, 3, and 10th eigenvalues of the observation Gramian.
@@ 3.4.3 Five true vortices and three-vortex estimator
In all other aspects, the problem remains unchanged from the previous example, except the initial true states are slightly different, given by $(x^{*}_{0}/R_{c},y^{*}_{0}/R_{c},\Gamma/UR_{c})=(-3,-0.8,-1)$, (2.8,0.5,1.1)(-2.8,-0.5,-1.1), (2.9,0,1.2)(-2.9,0,1.2), (2.7,0.5,1.3)(-2.7,0.5,1.3) and (2.8,0.8,1.4)(-2.8,0.8,1.4). Additionally, the dimension of the estimated state is now reduced to $n=9$, compared to 15 in the previous case. Figure 13(a) shows the true and estimated trajectories. Obviously the three estimated vortices cannot precisely follow the trajectories of the five true ones. However, they do appear to estimate the cluster trajectories well. Indeed, Fig. 13(b) confirms this clustering effect in the estimated vortex strengths: the estimated vortex nearest the two negatively signed true vortices below the cylinder is estimated to have a strength nearly equal to the sum of those two true vortices, denoted by $\Gamma_{1}$ and $\Gamma_{2}$. Similarly, the green and blue estimated vortices, which are positioned between the uppermost true vortices, $\Gamma_{3}$ through $\Gamma_{5}$, approximate the combined circulation of these three vortices.
@@
Figure 13: Five true vortices and three-vortex estimator. See Fig. 11 for legend.
@@ 3.5 Comments on sequential estimation
In Section 2 we stressed the importance of a good prior for successful flow estimation. In sequential estimation we achieve this by leveraging the knowledge obtained in previous steps. To underscore this, we return to the example of da Silva and Colonius [da2018ensemble], in which they sought to estimate the grid-based velocity field of flows past various bodies with a modest number of sensors. Without a good prior, it is hopeless for the measurements from a few velocity sensors to correct the high-dimensional grid velocity. However, the forecast model (the Navier–Stokes equations and boundary conditions) is deterministic, so even if the initial ensemble is drawn from a poor distribution—e.g., a spatially-uncorrelated Gaussian distribution, or white noise—the state’s ensemble is shaped over several forecast steps into a narrower and physically-plausible part of the state space. Indeed, in some of the examples in [da2018ensemble], the Navier-Stokes updates move the ensemble toward a trajectory on which the only uncertainty is in the phase of vortex shedding. Provided that there is sufficient initial variance of this phase among the ensemble members, the measurements are readily able to synchronize the evolving state onto the true phase. Even so, the benefits of this sequential improvement are still assisted by a good initial prior. Da Silva and Colonius [da2018ensemble] found the best performance when the initial states in their ensemble of $M$ members were drawn from the space spanned by the leading $M$ proper orthogonal decomposition (POD) modes of the true flow. In the vortex estimation examples in this section, our initial prior consisted of non-overlapping Gaussian distributions for each estimator vortex, each of which was reasonably close to a true vortex.
@@
The last example in Section 3.4, in which we interpreted measurements from a higher-dimensional system (five vortices) with a lower-dimensional representation (three vortices), provides an opportunity to make an important and more general point. The estimated trajectories in Fig. 13 show that the state covariance is quite small during most of the vortex–body interaction. However, it is crucial that we do not interpret this small covariance as precise knowledge of the true state. Rather, it represents a precise interpretation of the given measurements in the space in which we have confined our search. But that three-vortex space is a lower-dimensional manifold within the true space, and is thus inherently accompanied by a deterministic model error that we have no hope of estimating with our simplified state representation and observation model [cohn1997introduction]. Fluid dynamics is replete with approximate representations of the flow field, each of which has a companion model error due to the restriction onto the simpler manifold, e.g., truncation error in grid-based CFD representations, subgrid-scale error in turbulence models, residual errors in reduced-order modal decompositions. For example, in the aforementioned study of da Silva and Colonius [da2018ensemble], in which they assembled their initial flow states from random combinations of POD modes, they might have instead chosen to work exclusively in the state space of POD coefficients. In that same spirit, in recent work [mousavi2025sequentialestimationdisturbedaerodynamic], the authors of this paper have performed stochastic estimation in a latent space learned from a neural-network autoencoder. Some researchers have shown that it is possible to augment the state vector with the model error itself and, by virtue of a model that relates it to the measurements, estimate this error along with the simplified flow state [da2020flow].
@@
Finally, we note that the LREnKF has provided a principled way to reduce the rank of the filtering so that the effect of spurious errors is mitigated, but there is a variety of other methods that improve the ensemble Kalman filter and we have only briefly listed a few. We also note that one could generalize the linear Kalman update (32) to a nonlinear form, thereby relaxing the assumption of Gaussian distributions in the analysis step. Such is the goal of the nonlinear ensemble filter of Spantini et al. [spantini2022coupling]. Recall that the joint distribution $\pi(\mathsf{x},\mathsf{y})$ can be partitioned in two ways in equation (5): into the prior and likelihood, or into the marginal $\pi(\mathsf{y})$ and the conditional $\pi(\mathsf{x}|\mathsf{y})$ (i.e., the posterior, which we ultimately seek). Figure 1 illustrated these partitions for a linear observation operator, and we made use of the nice properties of the Gaussian to extract the posterior. However, another way to extract this posterior is to map the joint distribution (the ellipse in Figure 1) into a standard normal distribution $\mathcal{N}(0,I_{n+d})$ (a sphere of radius 1). The latter is trivial to partition into $\mathcal{N}(0,I_{d})\mathcal{N}(0,I_{n})$, and we can map the latter $n$-dimensional part back to the state space for the desired posterior. In fact, this invertible mapping is always possible for any distribution, and this is the basic idea used in [spantini2022coupling]. For this to be useful, we first have to have some knowledge of the joint distribution $\pi(\mathsf{x},\mathsf{y})$, and we can obtain such knowledge by assembling an ensemble of prior states with their associated predicted noisy measurements (8). The method has been preliminarily evaluated in an aerodynamics problem by Le Provost et al. [le2021low], but further investigation is needed to explore its merits in this context.
@@ 4 Neural networks and other learned mappings
Deep neural networks (NN) and other tools from machine learning have provided a fertile ground in recent years for learning operators that estimate flow quantities from aerodynamic measurements. Zhong et al. [zhong2023sparse] have provided a pertinent example of this in the unsteady aerodynamics context. In that work, a small number of surface pressure measurements obtained from a CFD simulation of an airfoil in a disturbed flow were provided to a multi-layer perceptron (MLP) network. The network’s resulting vector was a high-dimensional representation of the flow field (though still lower dimensional than the flow itself). This vector was then used by a convolutional neural network (CNN) to reconstruct the vorticity field or by another MLP to estimate the instantaneous lift. Morra et al. [morra2024ml] have recently shown that a trained deep operator network can be used in place of the full compressible Navier–Stokes equations to map upstream perturbations to downstream surface pressure measurements in a high-speed transitional boundary layer on a cone. They successfully used this approximate operator to estimate the amplitude of an upstream disturbance from downstream measurements taken from an experiment of the same geometry. Our objective in this paper is not to review the tools from machine learning and their application to fluid dynamics. Rather, our goal is to examine how to use these tools to support the stochastic flow estimation tasks that form the theme of this paper.
Among current architectures of deep neural networks, the autoencoder provides the most powerful for distilling data into low-dimensional representations. During unsupervised training on a set of high-dimensional vectors, an autoencoder simultaneously learns to compress this data into a low-dimensional latent space and to expand latent vectors back to the high-dimensional space with minimal error. We specifically call attention to the lift-augmented autoencoder developed recently by Fukami and Taira [fukami2023grasping] (and illustrated in Figure 14), in which the addition of a NN to predict lift helps to regularize an autoencoder that learns to compress vorticity field snapshots from a wide variety of undisturbed and disturbed two-dimensional flows past an airfoil at several angles of attack. The resulting latent space is three-dimensional, i.e., only three components are needed to uniquely specify the instantaneous flow field!
Figure 14: Lift-augmented autoencoder for non-linear reduced order of flow states.
@@
As with physics-based operators, our overarching objective with learned operators is to understand how uncertainty in the measurements $\mathsf{y}$ translates into uncertainty in our state estimate $\mathsf{x}$—in this case, via uncertainty in the latent state $\boldsymbol{\xi}$. If we combine the NN operators with models for the measurement and process noise, e.g., zero-mean Gaussians, then we have the ingredients of the sequential filtering equations—the transition probability (27) and the likelihood (29)—and these enable us to track uncertainty through the flow prediction. This was the approach taken by Mousavi and Eldredge [mousavi2025sequentialestimationdisturbedaerodynamic], in which pressure measurements were used to estimate a seven-dimensional latent vector representation of the full flow state. The latent space and associated operators were learned from snapshots of the disturbed flow past a wing at several fixed angles of attack. But there are important questions to explore before pursuing this: How does uncertainty propagate through the latent space? And how does uncertainty in the model itself enter the problem? In this section we will explore these questions by focusing on the approximate inverse observation operator, $\tilde{g}(\mathsf{y};w)$, which maps measurements $\mathsf{y}$ to the latent state $\boldsymbol{\xi}$. We will use this opportunity to quantify the various forms of uncertainty in this mapping.
Figure 15: Neural network architecture for flow prediction from sparse noisy pressure measurements $\boldsymbol{p}$, adapted with permission from Mousavi and Eldredge (J. Fluid Mech., 2025) [mousavi2025loworderflowreconstructionuncertainty]. Baseline (undisturbed) trajectories (in color); disturbed trajectories (gray, with one case highlighted in black).
In our aerodynamics context, a representative example of this inverse mapping was investigated in [mousavi2025loworderflowreconstructionuncertainty], in which a lift-augmented autoencoder was trained on CFD simulation data from the same undisturbed and disturbed airfoil flows as in [fukami2023grasping]. After training the autoencoder, the three-dimensional latent vector $\boldsymbol{\xi}_{j}$ for each training snapshot $j$ was paired with a measurement vector, $\mathsf{y}_{j}$, of the corresponding CFD-computed pressures at eleven evenly-spaced sensors on the surface of the airfoil. This latter vector also included the sensors’ position coordinates to distinguish angles of attack and to respect the sensors’ relative positioning. We will denote the training set of all such data pairs by $D=\{\mathsf{y}_{j},\boldsymbol{\xi}_{j}\}_{j=1}^{M}$.
@@
However, because the neural network provides us with access to the Jacobian $\nabla_{\mathsf{y}}\tilde{g}$ (through automatic differentiation), we can use this to characterize the uncertainty propagation more accurately. One such analytical method, stable distribution propagation, was introduced by Petersen et al. [petersen2024uncertaintyquantificationstabledistribution] and is applicable to Gaussian and Cauchy-distributed uncertainties. The core idea is to approximate the network’s behavior via local linearization. In the context of $\tilde{g}$, the expected value and variance of the predicted output are estimated as